OpenStudy (anonymous):

What is height of the pole? A. 8 feet B. 10 feet C. 36 feet D. 48 feet

OpenStudy (anonymous):

here

OpenStudy (anonymous):

How would I solve it?

OpenStudy (anonymous):

similar triangles

OpenStudy (anonymous):

It looks like it would be D?

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\frac{x}{64}=\frac{36}{x}\] \[x^2=64\times 36\] \[x=\sqrt{64\times 36}=8\times 4=48\]

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

oops typo on last line \[x^2=\sqrt{64\times 36}=8\times 6=48\]

OpenStudy (anonymous):

answer was right, but i put 4 in instead of 6, that was a mistake yw